OpenStudy (anonymous):

What is the center and radius of the circle with equation x2 + y2 – 8x – 6y – 75 = 0?

OpenStudy (anonymous):

you gotta complete the square on x and y

OpenStudy (anonymous):

Oh ok I got you ty

OpenStudy (anonymous):

Would it be (4,3) ?

OpenStudy (mathteacher1729):

xJai, how did you obtain (4,3) ? Like, what specific steps did you go through...

OpenStudy (anonymous):

Combining the like terms halfing then squaring the terms with the x then rationalizing them then once rationalized you would make them equal 0

OpenStudy (mathteacher1729):

(4,3) is correct, and it sounds like you know how to complete the square. cheers. :)

OpenStudy (anonymous):

do you know how to find the radius xjai?

OpenStudy (anonymous):

(x-4)^2 -16 + (y-3)^2 - 9 = 75 (x-4(^2 + (y-3)^2 = 100 so radius is 10

OpenStudy (anonymous):

center (4,3) and radius 10